INAF queue observing, 2015 Jan 15-16

Observer: E. Sani, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Michelle Edwards (on call) and Barry Rothberg (at the telescope)
Instrument: MODS - LBCred

Summary

During the twilight the sky is mostly clear, but we have stron wind with maximum gusts at 24.5 m/s. We can't run skyflats with LBCred.
At the dome opening we have cirrus scattered over the south horizon, maximum wind gusts 17.5 m/s and low humidity (16%).
Extremely variable seeing during the night ranging from 0.7 to 1.8.

Data Summary

Project_RUN Objects Filter ExpTime Notes
SaraccoB RDCS red gray only filter r_SDSS 2h variable seeing
MIgnoli MOS22 red grat only filter r_SDSS 2 h pseudomonucular mode. variable seeing
MIgnoli LBC_red_MOS22 filter z_SDSS 30m  
MIgnoli LBC_red_MOS22 filter i_SDSS 30m  
MIgnoli LBC_red_MOS22 filter r_SDSS 30m  
MIgnoli LBC_red_MOS22 filter Y-FAN 12m  
MIgnoli LBC_red SA104 standard filter r_, i_, z_SDSS, Y-FAN   airmass 1.15

Details

07:21 The TO opens the dome

07:37 acqMODS mods.7.rec.acq mods1r.20150116.0002-0004
  • we need to unwrap and send again the acquisition

07:58 execMODS mods.7.RDCSm1_5.obs mods1r.0005-0008
  • the average seeing is 1.2 at the beginning of integration.
  • while integrating the 3rd spetrum the seeing increases: 1.4-1.5
  • even worse at the end of the integration: 1.8

09:08 execMODS mods.7.RDCSm1_6.obs mods1r.0009-0010
  • the average seeing is 1.6 at the beginning of integration.
  • during the first integration the seeing reaches 1.9, then goes down again to 1.4

09:52 We have strong spherical aberration. This is because the star is too close to the edge of the patrol field.
  • We stop the 3rd integration and change the guiding star. The telescope is still tracking
  • The new GS is at the edge of the patrol field, corresponds to an allignment star, and quite faint (R=16).
  • The sky is anyhow clear and the GCS is fine in collimating.
  • We are confident in the mask alignment thanks to the previous recoveries
  • The TO clear the optics and applies manual corrections because it takes a while to collimate
  • acqMODS mods.7.rec.acq mods1r.0011
  • too many aberrations, the TO shifts to a nearby star

10:10 We slew to Mignoli_37 MOS22 and collimate with a bright star
  • acqMODS mods.22.rec.acq mods1r.0012
  • run dofpia, /first, /x2, /redonly on the scientific field. Coma aberration. run it again
  • the recovery failed

10:25 acqMODS mods.22..acq mods1r.0013-00016

10:28 run 1_MOS22_z.ob
  • image lbcr.20150116.103245 and lbcr..104740 have a jump due to a mods offset

10:47 execMODS mods.22.2.obs mods1r.0017-0018
  • at the beginning of the integration the average seeing is 1.1

11:10 run 2_MOS22_i.ob
  • run dofpia, /redonly on the scientific field
  • pause it after the 3rd image because MODS was committing an offset

11:23 execMODS mods.22.3.obs mods1r.0019-0022
  • start again 2_MOS22_i.ob
  • jump in lbcr.112449

12:03 run 3_MOS22_r.ob
  • run dofpia, /redonly

12:34 execMODS mods.22.3.obs mods1r.0023-0024

12:48 run dofpia, /redonly on the scientific field
  • run 4_MOS22_y.ob.
  • stop the OB after 4th image
  • DIMM seeing alwais >1.1
01:13 Slew to the standard star
  • run SA104_rizY.ob
  • after the rSloan filter the LBC is no doing nothing.
  • airmass 1.15

01:43 Run sky flats only on the red channel
  • SkyFlat _rz_0.ob GOOD scaling factor 3
  • SkyFlat _rz_180.ob GOOD scaling factor 1
  • SkyFlat _ry_180.ob only 2 frames are good, scaling factor 0.5
  • Good flats in the i_Sloan filter and r_Sloan filter can be find in /Repository/20141229 and /20141224 respectively

 
Topic revision: r11 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback